# Infrared fixed point in the massless twelve-flavor SU(3) gauge-fermion system

### 28 Feb 2024 v1

@@ Abstract
We present strong numerical evidence for the existence of an infrared fixed point in the renormalization group flow of the SU(3) gauge-fermion system with twelve massless fermions in the fundamental representation. Our numerical simulations using nHYP-smeared staggered fermions with Pauli-Villars improvement do not exhibit any first-order bulk phase transition in the investigated parameter region. We utilize an infinite volume renormalization scheme based on the gradient flow transformation to determine the renormalization group $\beta$ function. We identify an infrared fixed point at $g^{2}_{\mathrm{GF}\star}=6.69(68)$ in the GF scheme and calculate the leading irrelevant critical exponent $\gamma_{g}^{\star}=0.204(36)$. Our prediction for $\gamma_{g}^{\star}$ is consistent with available literature at the $1\mbox{-}2\sigma$ level.
@@ I Introduction
The infrared properties of the SU(3) gauge-fermion system with $N_{f}=12$ massless fundamental flavors has been studied extensively using a variety of analytical and numerical techniques. Such techniques include perturbation theory Ryttov and Shrock (2011); Pica and Sannino (2011); Ryttov and Shrock (2016a, b); Di Pietro and Serone (2020), use of the gap equation Appelquist et al. (1998); Bashir et al. (2013), functional renormalization group methods Braun and Gies (2006); Braun et al. (2011), conformal expansion Lee (2021), conformal bootstrap Li and Poland (2020), the background field method Grable and Romatschke (2023), perturbative non-relativistic quantum chromodynamics Chung and Nogradi (2023), large-$N$ expansion Romatschke (2024), and nonperturbative lattice simulations Appelquist et al. (2008, 2009); Lin et al. (2012, 2015); Fodor et al. (2016); Hasenfratz and Schaich (2018); Fodor et al. (2018a, b); Hasenfratz et al. (2019a, b); Hasenfratz (2010, 2012); Hasenfratz and Witzel (2019); Deuzeman et al. (2010); Fodor et al. (2011); Appelquist et al. (2011); DeGrand (2011); Fodor et al. (2012a); Aoki et al. (2012, 2013); Cheng et al. (2014); Lombardo et al. (2014); Fodor et al. (2009); Cheng et al. (2013). Investigations based on lattice simulations have utilized finite-volume step-scaling Appelquist et al. (2008, 2009); Lin et al. (2012, 2015); Fodor et al. (2016); Hasenfratz and Schaich (2018); Fodor et al. (2018a, b); Hasenfratz et al. (2019a, b), Monte Carlo renormalization group methods Hasenfratz (2010, 2012), hadron mass and decay constant spectroscopy Deuzeman et al. (2010); Fodor et al. (2011); Appelquist et al. (2011); DeGrand (2011); Fodor et al. (2012a); Aoki et al. (2012, 2013); Cheng et al. (2014); Lombardo et al. (2014), and the Dirac eigenmode spectrum Fodor et al. (2009); Cheng et al. (2013). Many investigations suggest that the $N_{f}=12$ system is infrared conformal [^1] See Refs. Ryttov and Shrock (2011); Pica and Sannino (2011); Ryttov and Shrock (2016a); Di Pietro and Serone (2020); Appelquist et al. (1998); Bashir et al. (2013); Lee (2021); Li and Poland (2020); Appelquist et al. (2008, 2009); Lin et al. (2012); Hasenfratz and Schaich (2018); Hasenfratz et al. (2019a, b); Hasenfratz (2010, 2012); Deuzeman et al. (2010); Appelquist et al. (2011); DeGrand (2011); Aoki et al. (2012, 2013); Cheng et al. (2014); Lombardo et al. (2014); Cheng et al. (2013), though a minority of studies conclude that the system is confining with chiral symmetry breaking, or are inconclusive, as they find neither direct evidence of chiral symmetry breaking nor of an infrared fixed point [^2] See Refs. Grable and Romatschke (2023); Romatschke (2024); Fodor et al. (2011); Lin et al. (2015); Fodor et al. (2012a, 2016, 2019).
@@
Figure 1: Our continuum prediction for $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ as a function of $g^{2}_{\mathrm{GF}}$ (maroon band) juxtaposed against the 1- (dashed), 2- (dotted) and 3-loop (dashed-dotted) gradient flow $\beta$ function from perturbation theory Harlander and Neumann (2016). The width of the maroon band indicates the error. Also given is our prediction for the leading irrelevant critical exponent at the IRFP, $\gamma_{g}^{\star}=0.204(36)$.
@@
As a preview, we show our nonperturbative prediction for $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ as a function of $g^{2}_{\mathrm{GF}}$ in Fig. 1. The predicted $\beta$ function converges to the universal 1-/2-loop and 3-loop gradient flow perturbative $\beta$ functions at small $g^{2}_{\mathrm{GF}}$ Harlander and Neumann (2016). Around $g^{2}_{\mathrm{GF}\star}=6.69(68)$, the nonperturbative $\beta$ function unambiguously exhibits an infrared fixed point. From the slope of $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ at $g^{2}_{\mathrm{GF}\star}$, we calculate the leading irrelevant critical exponent $\gamma_{g}^{\star}=0.204(36)$ and find that it is consistent with the perturbative calculations of Refs. Ryttov and Shrock (2016b); Di Pietro and Serone (2020) at the $1\sigma$ level and the lattice calculation of Ref. Hasenfratz and Schaich (2018) at the $2\sigma$ level. We control for systematical errors in the infinite volume extrapolation step of the CBFM using Bayesian model averaging. Additionally, our Pauli-Villars improved simulations offer tight control over systematics in the continuum extrapolation step of the CBFM. Our data is publicly available at Ref. Peterson and Hasenfratz (2024).
@@ III.1 Infinite volume extrapolation
In Figs. 2 and 3, we show the result of our model-averaged infinite volume extrapolation for the W and C discretization of $E(t)$, respectively, over a range of flow times $2.5\leq t/a^{2}\leq 6.0$ (different colors). The left panels of Figs. 2-3 show our infinite extrapolation of $g^{2}_{\mathrm{GF}}\big{(}t;L,g_{0}^{2}\big{)}$, while the right panels show our infinite volume extrapolation of $\beta_{\mathrm{GF}}\big{(}t;L,g_{0}^{2}\big{)}$. The bare gauge couplings that we chose for these plots are in the vicinity where the continuum $\beta$ function predicts an IRFP. For all three bare gauge couplings shown in Figs. 2-3, the model average is dominated by subsets containing $L/a\in\{28,32,36,40\}$, as $L/a=24$ often deviates from the linear trend in $a^{4}/L^{4}$, particularly as the flow time increases. This is reflected in the model average, as fits including $L/a=24$ possess small model weights $w_{\eta}$ and contribute negligibly to the model average.
@@ III.2 Intermediate interpolation
At each $t/a^{2}$, we account for the uncertainty in $g^{2}_{\mathrm{GF}}\big{(}t;g_{0}^{2}\big{)}$ by including the mean and covariance of $g^{2}_{\mathrm{GF}}\big{(}t;g_{0}^{2}\big{)}$ as a Gaussian prior. We also set a Gaussian prior on each coefficient $p_{n}$ with zero mean and a width of $0.1$, which helps stabilize the fit. We choose $N=4$, as it is the lowest value of $N$ that fits the data well. In Fig. 4, we show the result of our interpolation for the W operator (top panel) and the C operator (bottom panel) for several flow time values in the range $2.5\leq t/a^{2}\leq 6.0$ (different colors). Our fits have p-values in the $35\%-100\%$ range, but they strongly skew toward the higher end. This could indicate that we are either overfitting or the errors in our data are overestimated. Reducing the order makes each interpolation significantly worse, as interpolations with $N\leq 3$ are unable to accommodate the varying curvature at weak/strong coupling. Therefore, use $N=4$ for our central analysis. We will discuss the systematic effect that is associated with the order $N$ in our estimate of $g^{2}_{\mathrm{GF}\star}$ and $\gamma_{g}^{\star}$ in Sec. IV.
@@ III.3 Continuum extrapolation
We show our prediction for the continuum $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ in Fig. 6 from the WW (gold band) and WC (maroon band) combination. The continuum predictions from both operators are consistent with one another across the entire range of investigated renormalized couplings $g^{2}_{\mathrm{GF}}$. At small $g^{2}_{\mathrm{GF}}$, the continuum $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ appears to converge to the 1-, 2-, and 3-loop perturbative gradient flow $\beta$ function Harlander and Neumann (2016). At $g^{2}_{\mathrm{GF}\star}\approx 6.69$, our continuum $\beta$ function predicts an infrared fixed point. The location of the fixed point is slightly below the predicted IRFP from the step-scaling calculation of Ref. Hasenfratz and Schaich (2018). Note that, because the calculation in Ref. Hasenfratz and Schaich (2018) was done in a different gradient-flow-based renormalization scheme, the predicted $g^{2}_{\mathrm{GF}\star}$ values do not have to agree. Our final result for $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ from both operators is provided as an ASCII file.
@@ IV The IRFP and its leading irrelevant critical exponent
Figure 8: Comparison of our value for $\gamma_{g}^{\star}=0.204(36)$ (maroon errorbar) against Ref. Hasenfratz and Schaich (2018) (teal error bar) and Ref. Di Pietro and Serone (2020) (dark gold error bar). The smaller error bar on our result indicates the error without accounting for systematic effects; the larger error bar indicates our error after accounting for systematic effects. We indicate our total error with a grey band for visualization.
@@
Steps (1)-(5) yield $g^{2}_{\mathrm{GF}\star}=6.77(39),6.69(40)$ and $\gamma_{g}^{\star}=0.210(22),0.204(21)$ for WW and WC, respectively. In Fig. 7 we look at how the continuum limit predictions for $g^{2}_{\mathrm{GF}\star}$ (top panels) and $\gamma_{g}^{\star}$ (bottom panels) vary with our choice of $t_{\mathrm{min}}/a^{2}$ (x-axes) and $t_{\mathrm{max}.}/a^{2}$ (different colors) for the W (left panels) and C (right panels) operators. The central values for both quantities and both operators are stable; they vary well within error. The stability in $t_{\mathrm{min}}/a^{2}$ is attributed to the linearity of the continuum extrapolation over a wide range of $a^{2}/t$, while the stability in $t_{\mathrm{max}}/a^{2}$ is likely attributed to our control over the infinite volume extrapolation.
@@
  • Choosing a higher-order polynomial for the intermediate interpolation in Sec. III.2. The highest-order interpolation in $N$ that we can use before we lose control over our continuum extrapolation due to overfitting is $N=7$. This shifts the value of $\gamma_{g}^{\star}$ by $\approx 0.003$, and we take the latter difference as an estimate for the systematic error that is associated with our choice of $N$ for the intermediate interpolation.
@@
  • Choosing instead the prediction for $\gamma_{g}^{\star}=0.210(22)$ from the WW combination as our central result. We take the difference in these predictions ($\approx 0.006$) as an estimate of the systematic error associated with making a particular choice of flow/operator combination.
To be conservative, we combine the error in our analysis of $\gamma_{g}^{\star}$ with the systematic error estimates above linearly. This yields the final of prediction $\gamma_{g}^{\star}=0.204(36)$. Repeating the same exercise for $g^{2}_{\mathrm{GF}\star}$ yields a systematic error of $\approx 0.11$ from the interpolation order, $\approx 0.09$ from the continuum extrapolation, and $\approx 0.08$ from the flow/operator combination. Including the systematic error in $g^{2}_{\mathrm{GF}\star}$ linearly yields a final prediction of $g^{2}_{\mathrm{GF}\star}=6.69(68)$.
In Fig. 8 we compare our prediction for $\gamma_{g}^{\star}$ with those available in the literature Hasenfratz and Schaich (2018); Di Pietro and Serone (2020). Our result is plotted as a maroon star with errors indicated by an error bar. The smaller error bar is our error estimate before accounting for systematic effects and the larger error bar includes systematic effects. The result for $\gamma_{g}^{\star}$ from the perturbative calculation of Ref. Di Pietro and Serone (2020) (dark gold error bar) is within $1\sigma$ of our estimate for $\gamma_{g}^{*}$. The lattice calculation of $\gamma_{g}^{\star}$ from Ref. Hasenfratz and Schaich (2018) (cyan error bar) is within $2\sigma$ of our result. However, it is important to note that the lattice calculation in Ref. Hasenfratz and Schaich (2018) uses smaller volumes and very coarse lattices in comparison to the present work. It is also worth noting the “scheme-independent” prediction of $\gamma_{g}^{\star}\approx 0.228$ from Ref. Ryttov and Shrock (2016b) is also within $1\sigma$ of our predicted $\gamma_{g}^{\star}$; we don’t show this result in Fig. 8 because no estimate of the systematic error in this result is available.

### 13 Mar 2024 v2

@@ Abstract
We present strong numerical evidence for the existence of an infrared fixed point in the renormalization group flow of the SU(3) gauge-fermion system with twelve massless fermions in the fundamental representation. Our numerical simulations using nHYP-smeared staggered fermions with Pauli-Villars improvement do not exhibit any first-order bulk phase transition in the investigated parameter region. We utilize an infinite volume renormalization scheme based on the gradient flow transformation to determine the renormalization group $\beta$ function. We identify an infrared fixed point at $g^{2}_{\mathrm{GF}\star}=6.60(62)$ in the GF scheme and calculate the leading irrelevant critical exponent $\gamma_{g}^{\star}=0.199(32)$. Our prediction for $\gamma_{g}^{\star}$ is consistent with available literature at the $1\mbox{-}2\sigma$ level.
@@ I Introduction
The infrared properties of the SU(3) gauge-fermion system with $N_{f}=12$ massless fundamental flavors has been studied extensively using a variety of analytical and numerical techniques. Such techniques include perturbation theory Ryttov and Shrock (2011); Pica and Sannino (2011); Ryttov and Shrock (2016a, 2017, b); Di Pietro and Serone (2020), use of the gap equation Appelquist et al. (1998); Bashir et al. (2013), functional renormalization group methods Braun and Gies (2006); Braun et al. (2011), conformal expansion Lee (2021), conformal bootstrap Li and Poland (2020), the background field method Grable and Romatschke (2023), perturbative non-relativistic quantum chromodynamics Chung and Nogradi (2023), large-$N$ expansion Romatschke (2024), and nonperturbative lattice simulations Appelquist et al. (2008, 2009); Lin et al. (2012, 2015); Fodor et al. (2016); Hasenfratz and Schaich (2018); Fodor et al. (2018a, b); Hasenfratz et al. (2019a, b); Hasenfratz (2010, 2012); Hasenfratz and Witzel (2019); Deuzeman et al. (2010); Fodor et al. (2011); Appelquist et al. (2011); DeGrand (2011); Fodor et al. (2012a); Aoki et al. (2012, 2013); Cheng et al. (2014); Lombardo et al. (2014); Fodor et al. (2009); Cheng et al. (2013). Investigations based on lattice simulations have utilized finite-volume step-scaling Appelquist et al. (2008, 2009); Lin et al. (2012, 2015); Fodor et al. (2016); Hasenfratz and Schaich (2018); Fodor et al. (2018a, b); Hasenfratz et al. (2019a, b), Monte Carlo renormalization group methods Hasenfratz (2010, 2012), hadron mass and decay constant spectroscopy Deuzeman et al. (2010); Fodor et al. (2011); Appelquist et al. (2011); DeGrand (2011); Fodor et al. (2012a); Aoki et al. (2012, 2013); Cheng et al. (2014); Lombardo et al. (2014), and the Dirac eigenmode spectrum Fodor et al. (2009); Cheng et al. (2013). Many investigations suggest that the $N_{f}=12$ system is infrared conformal [^1] See Refs. Ryttov and Shrock (2011); Pica and Sannino (2011); Ryttov and Shrock (2016a, 2017); Di Pietro and Serone (2020); Appelquist et al. (1998); Bashir et al. (2013); Lee (2021); Li and Poland (2020); Appelquist et al. (2008, 2009); Lin et al. (2012); Hasenfratz and Schaich (2018); Hasenfratz et al. (2019a, b); Hasenfratz (2010, 2012); Deuzeman et al. (2010); Appelquist et al. (2011); DeGrand (2011); Aoki et al. (2012, 2013); Cheng et al. (2014); Lombardo et al. (2014); Cheng et al. (2013), though a minority of studies conclude that the system is confining with chiral symmetry breaking, or are inconclusive, as they find neither direct evidence of chiral symmetry breaking nor of an infrared fixed point [^2] See Refs. Grable and Romatschke (2023); Romatschke (2024); Fodor et al. (2011); Lin et al. (2015); Fodor et al. (2012a, 2016, 2019).
@@
Figure 1: Our continuum prediction for $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ as a function of $g^{2}_{\mathrm{GF}}$ (maroon band) juxtaposed against the 1- (dashed), 2- (dotted) and 3-loop (dashed-dotted) gradient flow $\beta$ function from perturbation theory Harlander and Neumann (2016). The width of the maroon band indicates the error. Also given is our prediction for the leading irrelevant critical exponent at the IRFP, $\gamma_{g}^{\star}=0.199(32)$.
@@
As a preview, we show our nonperturbative prediction for $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ as a function of $g^{2}_{\mathrm{GF}}$ in Fig. 1. The predicted $\beta$ function converges to the universal 1-/2-loop and 3-loop gradient flow perturbative $\beta$ functions at small $g^{2}_{\mathrm{GF}}$ Harlander and Neumann (2016). Around $g^{2}_{\mathrm{GF}\star}=6.60(62)$, the nonperturbative $\beta$ function unambiguously exhibits an infrared fixed point. From the slope of $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ at $g^{2}_{\mathrm{GF}\star}$, we calculate the leading irrelevant critical exponent $\gamma_{g}^{\star}=0.199(32)$ and find that it is consistent with the perturbative calculations of Refs. Ryttov and Shrock (2017); Di Pietro and Serone (2020) at the $1\sigma$ level and the lattice calculation of Ref. Hasenfratz and Schaich (2018) at the $2\sigma$ level. We control for systematic errors in the infinite volume extrapolation step of the CBFM using Bayesian model averaging. Additionally, our Pauli-Villars improved simulations offer tight control over systematics in the continuum extrapolation step of the CBFM. Our data is publicly available at Ref. Peterson and Hasenfratz (2024).
@@ II Numerical details
Figure 2: The gradient-flowed Polyakov loop expectation value at flow time $8t/a^{2}\approx(L/2a)^{2}$ versus the bare gauge coupling $\beta_{b}$ on each volume in Table 1. The absolute value of the Polyakov loop is shown by colored error bars: $L/a=24$ (blue), $28$ (yellow), $32$ (green), $36$ (orange), and $40$ (pink).
@@
In Fig. 2 we plot the expectation value of the magnitude of the gradient-flowed Polyakov loop $|P|$ at $8t/a^{2}\approx(L/2a)^{2}$ against the bare gauge coupling $\beta_{b}$ for all ensembles in Table. 1. Errors are estimated using the “$\Gamma$-method” technique implemented by the pyerrors package Wolff (2004); Schaefer et al. (2011); Ramos (2021); Joswig et al. (2023a, b); however, they are likely underestimated. Nonetheless, the Polyakov loop suggests that the system is not confining in the range of couplings and volumes that we use in the present study.
@@ III.1 Infinite volume extrapolation
In Figs. 3 and 4, we show the result of our model-averaged infinite volume extrapolation for the W and C discretization of $E(t)$, respectively, over a range of flow times $2.5\leq t/a^{2}\leq 6.0$ (different colors). The left panels of Figs. 3-4 show our infinite extrapolation of $g^{2}_{\mathrm{GF}}\big{(}t;L,g_{0}^{2}\big{)}$, while the right panels show our infinite volume extrapolation of $\beta_{\mathrm{GF}}\big{(}t;L,g_{0}^{2}\big{)}$. The bare gauge couplings that we chose for these plots are in the vicinity where the continuum $\beta$ function predicts an IRFP. For all three bare gauge couplings shown in Figs. 3-4, the model average is dominated by subsets containing $L/a\in\{28,32,36,40\}$, as $L/a=24$ often deviates from the linear trend in $a^{4}/L^{4}$, particularly as the flow time increases. This is reflected in the model average, as fits including $L/a=24$ possess small model weights $w_{\eta}$ and contribute negligibly to the model average.
@@ III.2 Intermediate interpolation
At each $t/a^{2}$, we account for the uncertainty in $g^{2}_{\mathrm{GF}}\big{(}t;g_{0}^{2}\big{)}$ by including the mean and covariance of $g^{2}_{\mathrm{GF}}\big{(}t;g_{0}^{2}\big{)}$ as a Gaussian prior. We also set a Gaussian prior on each coefficient $p_{n}$ with zero mean and a width of $0.1$, which helps stabilize the fit. We choose $N=4$, as it is the lowest value of $N$ that fits the data well. In Fig. 5, we show the result of our interpolation for the W operator (top panel) and the C operator (bottom panel) for several flow time values in the range $2.5\leq t/a^{2}\leq 6.0$ (different colors). The fits that enter our continuum extrapolation have p-values in the $83\%-98\%$ range. This could indicate that we are either overfitting or the errors in our data are overestimated. Reducing the order makes each interpolation significantly worse, as interpolations with $N\leq 3$ are unable to accommodate the varying curvature at weak/strong coupling. Therefore, use $N=4$ for our central analysis. We will discuss the systematic effect that is associated with the order $N$ in our estimate of $g^{2}_{\mathrm{GF}\star}$ and $\gamma_{g}^{\star}$ in Sec. IV.
@@ III.3 Continuum extrapolation
We show our prediction for the continuum $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ in Fig. 7 from the WW (gold band) and WC (maroon band) combination. The continuum predictions from both operators are consistent with one another across the entire range of investigated renormalized couplings $g^{2}_{\mathrm{GF}}$. At small $g^{2}_{\mathrm{GF}}$, the continuum $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ appears to converge to the 1-, 2-, and 3-loop perturbative gradient flow $\beta$ function Harlander and Neumann (2016). At $g^{2}_{\mathrm{GF}\star}\approx 6.60$, our continuum $\beta$ function predicts an infrared fixed point. The location of the fixed point is slightly below the predicted IRFP from the step-scaling calculation of Ref. Hasenfratz and Schaich (2018). Note that, because the calculation in Ref. Hasenfratz and Schaich (2018) was done in a different gradient-flow-based renormalization scheme, the predicted $g^{2}_{\mathrm{GF}\star}$ values do not have to agree. Our final result for $\beta_{\mathrm{GF}}\big{(}g^{2}_{\mathrm{GF}}\big{)}$ from both operators is provided as an ASCII file.
@@ IV The IRFP and its leading irrelevant critical exponent
Figure 9: Comparison of our value for $\gamma_{g}^{\star}$ (maroon errorbar) against Ref. Hasenfratz and Schaich (2018) (teal error bar) and Ref. Di Pietro and Serone (2020) (dark gold error bar). The smaller error bar on our result indicates the error without accounting for systematic effects; the larger error bar indicates our error after accounting for systematic effects. We indicate our total error with a grey band for visualization.
@@
Steps (1)-(5) yield $g^{2}_{\mathrm{GF}\star}=6.69(35),6.60(36)$ and $\gamma_{g}^{\star}=0.206(19),0.199(18)$ for WW and WC, respectively. In Fig. 8 we look at how the continuum limit predictions for $g^{2}_{\mathrm{GF}\star}$ (top panels) and $\gamma_{g}^{\star}$ (bottom panels) vary with our choice of $t_{\mathrm{min}}/a^{2}$ (x-axes) and $t_{\mathrm{max}.}/a^{2}$ (different colors) for the W (left panels) and C (right panels) operators. The central values for both quantities and both operators are stable; they vary well within error. The stability in $t_{\mathrm{min}}/a^{2}$ is attributed to the linearity of the continuum extrapolation over a wide range of $a^{2}/t$, while the stability in $t_{\mathrm{max}}/a^{2}$ is likely attributed to our control over the infinite volume extrapolation.
@@
  • Choosing a higher-order polynomial for the intermediate interpolation in Sec. III.2. The highest-order interpolation in $N$ that we can use before we lose control over our continuum extrapolation due to overfitting is $N=6$. This shifts the value of $\gamma_{g}^{\star}$ by $\approx 0.001$, and we take the latter difference as an estimate for the systematic error that is associated with our choice of $N$ for the intermediate interpolation.
@@
  • Choosing instead the prediction for $\gamma_{g}^{\star}=0.206(19)$ from the WW combination as our central result. We take the difference in these predictions ($\approx 0.007$) as an estimate of the systematic error associated with making a particular choice of flow/operator combination.
To be conservative, we combine the error in our analysis of $\gamma_{g}^{\star}$ with the systematic error estimates above linearly. This yields the final of prediction $\gamma_{g}^{\star}=0.199(32)$. Repeating the same exercise for $g^{2}_{\mathrm{GF}\star}$ yields a systematic error of $\approx 0.12$ from the interpolation order, $\approx 0.05$ from the continuum extrapolation, and $\approx 0.09$ from the flow/operator combination. Including the systematic error in $g^{2}_{\mathrm{GF}\star}$ linearly yields a final prediction of $g^{2}_{\mathrm{GF}\star}=6.60(62)$.
In Fig. 9 we compare our prediction for $\gamma_{g}^{\star}$ with those available in the literature Hasenfratz and Schaich (2018); Di Pietro and Serone (2020). Our result is plotted as a maroon star with errors indicated by an error bar. The smaller error bar is our error estimate before accounting for systematic effects and the larger error bar includes systematic effects. The result for $\gamma_{g}^{\star}$ from the perturbative calculation of Ref. Di Pietro and Serone (2020) (dark gold error bar) is within $1\sigma$ of our estimate for $\gamma_{g}^{*}$. The lattice calculation of $\gamma_{g}^{\star}$ from Ref. Hasenfratz and Schaich (2018) (cyan error bar) is within $2\sigma$ of our result. However, it is important to note that the lattice calculation in Ref. Hasenfratz and Schaich (2018) uses smaller volumes and very coarse lattices in comparison to the present work. It is also worth noting the “scheme-independent” prediction of $\gamma_{g}^{\star}\approx 0.228$ from Ref. Ryttov and Shrock (2017) is also within $1\sigma$ of our predicted $\gamma_{g}^{\star}$; we don’t show this result in Fig. 9 because no estimate of the systematic error in this result is available.